• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 15:00:56 on April10,2008

    Entry number 231215

    keyword=L-HRS dipole problem/solution

    When we changed to p0 = 1.007 GeV on L-HRS, the dipole NMR would not
    lock. We power-cycled hareboot17 but this probably was not necessary.
    Called Javier Gomez. He found that probe A was locking at 0.17 T
    and tried to bump up the DAC but this didn't work. He switched to
    probe B and it worked -- found resonance. The dipole is settling
    in towards p0 = 1.007 now.